Respondents Sex Statistic Std. Error 6.33 5.12 7.54 4.85 2.00 90.364 9.506 EmailHrsiperWeek Male Mean 95% Confidence Interval for Mean 5% Trimmed Mean Median Variance Std. Deviation Minimum Maximum Range Interquartile Range Skewness Kurtosis 615 Lower Bound Upper Bound 50 50 2.658 7.585 5.93 4.98 6.88 4.57 2.00 78.924 8.884 157 314 485 Female Mean 95% Confidence Interval for Mean 5% Trimmed Mean Median Variance Std. Deviation Minimum Maximum Range Interquartile Range Skewness Kurtosis Lower Bound Upper Bound 50 50 2.675 7.834 133 265

Gender Male Female Variables EMAILHR EMAILHR Median 2.00 2.00 First quartile Third quartile
Answer #1

male

first quartile is the mid value between minimum value and median

minimum =0, median = 2

therefore first quartile = (2-0)/2= 1

third quartile is the mid value between median and maximum value

median = 2, maximum = 50

therefore, third quartile = (50-2)/2= 24

female

first qurtile = 1

third quartile = 24